OpenStudy (anonymous):

simplify. √75 -4√8 -3√32 ? a. -15√5 b. -15√6 c. 5√3 -20√2 d. 5√3 + 4√2

OpenStudy (anonymous):

root75 =root(25*3)=5* sqrt3 sqrt8=sqrt(4*2)=2 sqrt2 sqrt32=sqrt(16*2)=4sqrt2

OpenStudy (anonymous):

5* sqrt3-4*2 sqrt2-3*4sqrt2 =5 sqrt3-8sqrt2-12sqrt2 =5sqrt3-20sqrt2

OpenStudy (anonymous):

thanks c:

OpenStudy (anonymous):

yeah c

OpenStudy (anonymous):

\[\huge \sqrt{75}-4\sqrt8-3\sqrt{32}\] \[\huge \rightarrow sqrt {25 \times 3}-4\sqrt{4 \times 2}-3\sqrt{16 \times 2}\] \[\huge \rightarrow \sqrt{5^2 \times 3}-4\sqrt{2^2 \times 2}-3\sqrt{4^2 \times 2}\] \[\huge \rightarrow 5\sqrt{3}-4\times 2\sqrt{2}-3 \times 4 \sqrt{ 2}\] \[\huge \rightarrow 5\sqrt{3}-8\sqrt{2}-12 \sqrt{ 2}\] \[\huge \rightarrow 5\sqrt{3}-20\sqrt{2}\]